# Spin Seebeck Effect as a Probe for Majorana Fermions in Kitaev Spin Liquids

### v1

@@ IV.2 Magnetic field dependence of tunnel spin current
Since the positive sign of the spin Seebeck effect is common between the FM KSL and the FFM state, a way to distinguish between the two is needed. One is the field dependence discussed above: $\tilde{J}_{S}$ increases with $B$ in the low-field KSL, but it decreases in the high-field FFM state. Another efficient way is the dependence on the magnetic field direction. Figure 10 shows the field angle dependence of the tunnel spin currents in the $ac$ plane (see Fig. 1): (a) The results for the low-field KSL obtained by the perturbation analysis and (b) those for the high-field FFM state by the linear spin-wave theory. In the perturbation theory, the tunnel spin current $\tilde{j}_{S}$ vanishes for ${\bf B}\parallel z$. This is because $s^{\pm(1)}$, $s^{\pm(2)}_{1}$, and $s^{\pm(2)}_{2}$ in Eqs. (10)–(12) are all zero for ${\bf B}\parallel z$, as $B^{x}=B^{y}=0$ and $A^{z}=0$. The same applies for ${\bf B}\parallel x$ and ${\bf B}\parallel y$. Thus, the tunnel spin current for the low-field KSL is minimized for ${\bf B}\parallel z$, while maximized for ${\bf B}\parallel c$. The other maxima with the same value as ${\bf B}\parallel c$ are found at $\theta\approx 20^{\circ}$ and $200^{\circ}$ due to the symmetry of $\pi$ rotation about the $[110]$ axis ($\theta\approx 54.7^{\circ}$). Meanwhile, the linear spin-wave theory for the FFM state predicts an opposite tendency: The tunnel spin current is maximized for ${\bf B}\parallel z$, while minimized for ${\bf B}\parallel c$, as shown in Fig. 10(b). This is due to the logarithmic divergence discussed in Sec. III.3. We note that the TDVP simulation data presented in Figs. 4 and 5, however, are not sufficient to discuss these issues presumably due to the limited system sizes. Further extensions are needed for more precise predictions for experiments.
@@ V Summary
In summary, we have studied the spin Seebeck effect in both the ferromagnetic and antiferromagnetic Kitaev spin liquids by using the real-time dynamics simulation, the perturbation analysis, and the linear spin-wave theory, based on the tunnel spin-current theory. There are two main discoveries. The first one is that the spin Seebeck effect is induced in the Kitaev spin liquid by the low-energy fractional Majorana quasiparticles that have no spin angular momentum. This is a qualitatively different phenomenon from the spin Seebeck effect in the one-dimensional quantum spin liquid induced by the fractional spinon excitations with nonzero spin angular momentum. The second discovery is that the spin current mediated by elementary excitations in quantum spin liquids can be positive or negative; namely, we found the sign change of the spin current in the Kitaev spin liquid between the ferromagnetic and antiferromagnetic Kitaev interactions. Our finding suggests the Majorana fermions contribute to the spin current with up- and down-spin like nature in the ferromagnetic and antiferromagnetic Kitaev spin liquid, respectively. This finding refutes a possibility that the spin current carried by the fractional excitations is always negative in the quantum spin liquids.
[…]
@@ Acknowledgements.
The authors would like to thank K. Fukui, H.-C. Jiang, K. Kobayashi, E. Saitoh, S. Trebst, and A. Tsukazaki for fruitful discussions. This work was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Numbers JP19H05825, JP20H00122, JP20H01830, JP20H01849, JP22H01179, JP22K03509, and JP23H03818, JSPS Grant-in-Aid for Scientific Research on Innovative Areas Grants Numbers JP19H05825, JP22H04480, JP22H05131 and JP23H04576, JST COI-NEXT Program Grant Number JPMJPF2221, and JST, CREST Grant Number JP-MJCR18T2, Japan. This work was also supported by the National Natural Science Foundation of China (Grant No. 12150610462). Numerical calculations were performed using the facilities of the Supercomputer Center, The Institute for Solid State Physics, The University of Tokyo.
@@ Appendix A Spin current in the ferromagnetic Heisenberg model
In this Appendix, we show how to compute the tunnel spin current $\tilde{J}_{S}$ in Eq. (6) for the FM Heisenberg model [67, 58, 71] shown in Fig. 4(c). The Hamiltonian reads
@@
where the sum of $\langle j,j^{\prime}\rangle$ is taken for all the nearest-neighbor spins on the honeycomb lattice; we take $J=1$. Without loss of generality, we consider ${\bf B}\parallel z$, for which all the spins are fully polarized to the $z$ direction in the ground state.
We calculate $\tilde{J}_{S}$ for the fully-polarized (FP) state by using the linear spin-wave theory. By the Holstein-Primakoff transformation, $S^{+}_{j}=b_{j}$, $S^{-}_{j}=b^{\dagger}_{j}$, and $S^{z}_{j}=\frac{1}{2}-n_{j}$ with $n_{j}=b^{\dagger}_{j}b_{j}$, the Hamiltonian is written by the boson operators $b_{j}$ and $b^{\dagger}_{j}$ as
@@
$\mathcal{H}^{\rm LSW}_{{\rm H}}=$ $-\sum_{\bf r}\left(g_{\bf q}b^{\dagger}_{{\bf q}\rm B}b_{{\bf q}\rm A}+{\rm h.c.}\right)$
@@
where $g_{\bf q}=(J/2)(e^{i{\bf q}\cdot{\bf a}_{1}}+e^{i{\bf q}\cdot{\bf a}_{2}}+1)$ with two translation vectors ${\bf a}_{1}$ and ${\bf a}_{2}$, and the index $\alpha$ denotes the sublattice A and B. This is easily diagonalized as
$\mathcal{H}^{\rm LSW}_{{\rm H}}=$ $\sum_{{\bf q}}\sum_{\sigma=\pm}\varepsilon_{{\bf q}\sigma}\eta^{\dagger}_{{\bf q}\sigma}\eta_{{\bf q}\sigma},$ (20)
@@
$X_{+}(\omega)$ $=\frac{\delta}{2}\sum_{\ell}\frac{|\langle\ell|b^{\dagger}_{j_{0}}|{\rm FP}\rangle|^{2}}{(E_{\ell}-E_{\rm FP}-\omega)^{2}+\delta^{2}},$ (23)
$=\frac{\delta}{2N_{\rm u}}\sum_{{\bf q},\sigma}\frac{1}{(\varepsilon_{{\bf q}\sigma}-\omega)^{2}+\delta^{2}}$ (24)
@@ Appendix B Matrix elements of $s^{\pm(1)}$ and $s^{\pm(2)}_{1}$
where the sum runs over half of the first Brillouin zone with $q_{x}>0$ and $\alpha=A$ or B, Eq. (8) is rewritten as
$\mathcal{H}_{0}=$ ${\sum_{\bf q}}^{\prime}{\boldsymbol{\gamma}}^{\dagger}_{\bf q}\mathsf{H}_{\bf q}{\boldsymbol{\gamma}}_{\bf q},$ (28)
@@
$\mathsf{U}_{\bf q}=\frac{1}{\sqrt{2}|f_{\bf q}|}\left(\begin{matrix}|f_{\bf q}|&if_{\bf q}\\ -if^{*}_{\bf q}&-|f_{\bf q}|\end{matrix}\right),$ (34)
@@
$\mathcal{H}_{0}=$ ${\sum_{\bf q}}^{\prime}\sum_{\sigma=\pm}\sigma|f_{\bf q}|\alpha^{\dagger}_{{\bf q}\sigma}\alpha_{{\bf q}\sigma},$ (35)
where ${\boldsymbol{\alpha}}^{\dagger}_{\bf q}=(\alpha^{\dagger}_{{\bf q}+}~{}\alpha^{\dagger}_{{\bf q}-})$, and ${\boldsymbol{\alpha}}^{\dagger}_{\bf q}={\boldsymbol{\gamma}}^{\dagger}_{\bf q}\mathsf{U}_{\bf q}$.
@@
Note that the energy of $|\ell\rangle=|{\bf q}-,{\bf q}^{\prime}-\rangle$ and $|\ell^{\prime}\rangle=|{\bf q}+,{\bf q}^{\prime}+\rangle$ are the same, since $E_{\ell}=E_{\ell^{\prime}}=|f_{\bf q}|+|f_{{\bf q}^{\prime}}|$. For the particle-hole states, the matrix element is given by
@@
The matrix elements for $s^{-(2)}_{1}$ are obtained by replacing $A^{\gamma}$ by $(A^{\gamma})^{*}$. If the sign of $K$ is reversed, $f_{\bf q}$ becomes $-f_{\bf q}$, and then, $\langle{\bf q}\pm,{\bf q}^{\prime}\pm|s^{\pm(1)}|\emptyset\rangle$ and $\langle{\bf q}+,{\bf q}^{\prime}-|s_{1}^{\pm(2)}|\emptyset\rangle$ change their signs, while $\langle{\bf q}+,{\bf q}^{\prime}-|s^{\pm(1)}|\emptyset\rangle$ and $\langle{\bf q}\pm,{\bf q}^{\prime}\pm|s_{1}^{\pm(2)}|\emptyset\rangle$ are intact. Therefore, the sign reversal of $K$ leads $\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle\to-\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle$.

### v2

@@ IV.2 Magnetic field dependence of tunnel spin current
Since the positive sign of the SSE is common between the FM KSL and the FFM state, a way to distinguish between the two is needed. One is the field dependence discussed above: $\tilde{J}_{S}$ increases with $B$ in the low-field KSL, but it decreases in the high-field FFM state. Another efficient way is the dependence on the magnetic field direction. Figure 10 shows the field angle dependence of the tunnel spin currents in the $ac$ plane (see Fig. 1): (a) The results for the low-field KSL obtained by the perturbation analysis and (b) those for the high-field FFM state by the linear spin-wave theory. In the perturbation theory, the tunnel spin current $\tilde{j}_{S}$ vanishes for ${\bf B}\parallel z$. This is because $s^{\pm(1)}$, $s^{\pm(2)}_{1}$, and $s^{\pm(2)}_{2}$ in Eqs. (10)–(12) are all zero for ${\bf B}\parallel z$, as $B^{x}=B^{y}=0$ and $A^{z}=0$. The same applies for ${\bf B}\parallel x$ and ${\bf B}\parallel y$ by symmetry. Thus, the tunnel spin current for the low-field KSL is minimized for ${\bf B}\parallel z$, while maximized for ${\bf B}\parallel c$. The other maxima with the same value as ${\bf B}\parallel c$ are found at $\theta\approx 20^{\circ}$ and $200^{\circ}$ due to the symmetry of $\pi$ rotation about the $[110]$ axis ($\theta\approx 54.7^{\circ}$). Meanwhile, the linear spin-wave theory for the FFM state predicts an opposite tendency: The tunnel spin current is maximized for ${\bf B}\parallel z$, while minimized for ${\bf B}\parallel c$, as shown in Fig. 10(b). This is due to the logarithmic divergence discussed in Sec. III.3. We note that the TDVP simulation data presented in Figs. 4 and 5, however, are not sufficient to discuss these issues presumably due to the limited system sizes. Further extensions are needed for more precise predictions for experiments.
IV.3 Experimental implication
Our results of the SSE are expected to provide a strong insight into the identification of KSL in the candidate materials. As almost all the candidates show a magnetic order at low temperature at zero magnetic field due to parasitic non-Kitaev interactions such as the Heisenberg interaction, at least two scenarios have been proposed for the field dependence: One is that the system exhibits a direct transition from the low-field ordered phase to the FFM phase, and the other is that the system first exhibits a transition to an intermediate quantum spin liquid phase, and then at a higher field, it shows another transition or crossover to the FFM phase  (see, e.g., Fig. 16 in Ref. [28]). The inelastic neutron scattering (INS) measurements [95] as well as some of the thermal Hall conductivity measurements [33, 35, 34, 36] for the prime candidate $\alpha$-RuCl3 support the latter scenario, even though the evidence remains inconclusive. While our numerics has a limitation mainly because of the system size and the perturbation theory is limited to the weak-field region, we anticipate that our results contribute to understanding of the behavior under the magnetic field and the effect of the non-Kitaev interactions.
We emphasize that the field angle dependence of the SSE would be particularly important to distinguish the nature of quantum spin liquids. For instance, the RVB state, which is a prototypical example of quantum spin liquids, is expected not to exhibit nontrivial field-angle dependence unlike the KSL, since it is isotropic in spin space. Thus, the SSE measurement in the intermediate-field region of the Kitaev candidate materials would serve as a good probe for not only the existence but also the nature of the possible quantum spin liquid.
Besides the peculiar field-angle dependence, the sign change of the SSE signal would also be important in experiments for the Kitaev candidate materials. While the Kitaev interactions in the first-generation candidate materials, the Ir-based oxides $A_{2}$rO3 ($A$=Na, Li) and $\alpha$-RuCl3, are thought to be FM as theoretically predicted by Jackeli and Khaliullin [25], there have been theoretical proposals for new generation, such as cobaltites and $f$-electron compounds. In particular, it was predicted that the Kitaev interaction can be AFM in the $f$-electron systems [96, 97]. Hence, our finding of the sign dependence of the SSE would be a useful experimental tool to identify the sign of the Kitaev interaction in the future materials exploration.
The SSE measurement represents a distinctive experimental technique, having several advantages over other methods used to study the Kitaev magnets although it alone cannot fully unveil the nature of the KSL. The tunnel spin current is given by the frequency integral of the transverse component of the dynamical susceptibility, which is also measured by the INS using polarized neutrons [98, 99, 100, 101, 95, 102, 103, 104]. However, the SSE has advantages over the INS in several aspects: The INS requires a large-scale experimental facility and a huge amount of samples, and it is not suitable for some candidate materials when they include atoms that absorb neutrons (e.g., the Ir-based materials). In comparison, the SSE measurement can be done in much smaller-scale, table-top experiments, with a nano to micrometer scale sample, even in a monolayer. The electron spin resonance (ESR) is a method for detecting the spin dynamics at zero wavenumber in a certain frequency range, and is complementary to the SSE measurement, in which the integral over the frequency and wavenumber is obtained [105]. The nuclear magnetic resonance (NMR) measures electron spin dynamics in the low-frequency limit and is qualitatively different from the information obtained by the SSE measurement [106, 107]. The Raman scattering provides information over a wide frequency range, with the direction of polarization of the spin excitations depending on the polarization of light [108, 109, 110]. This situation is similar in the SSE measurement, in which the polarization direction is identified by the direction of the magnetic field. Nevertheless, we emphasize that the SSE has an advantage as it can detect the sign of the Kitaev interaction, which is one of the main results of this study. The thermal Hall effect and the SSE probe different nature of the Majorana excitations: The former observes their topological nature through the Berry curvature of the Majorana dispersions, while the latter observes their angular momenta. We anticipate that a comprehensive approach, combining results from multiple experimental methods including the SSE measurement, will shed further light on the intriguing phenomenon in the KSL.
IV.4 Experimental observability
In Sec. III.1, we discussed that the tunnel spin current for the KSL is observable in experiments from the comparison to the FM Heisenberg model on the two-dimensional honeycomb lattice. Here, we further reinforce the experimental observability by additional comparison with a 3D model in a realistic experimental setup as done in the previous studies [72, 62, 76]. Specifically, we consider the $S=1/2$ Heisenberg model with the nearest-neighbor FM interaction $J$ under an external magnetic field $B$ on a simple cubic lattice, and calculate the tunnel spin current $\tilde{J}_{S}$ by the spin-wave theory in Appendix A. To evaluate the magnitude of $\tilde{J}_{S}$, we choose typical values of $B$ and $T$ that facilitate the observation in experiments. Table 1 summarizes the calculated $\tilde{J}_{S}$ in the 3D FM state with $B/J=0.1$, $T/J=0.1$, and $0.5$, and those in the FM and AFM KSLs with $T=0.05$ [Fig. 4(c)]. Note that the critical temperature for the 3D FM state at $B=0$ is $T_{C}=0.839(1)$ [111, 112]. These results suggest that the signal intensity in the KSLs is expected to be $\sim 10^{-2}$$10^{0}$ of those in the typical 3D FM state. Compared to the signal intensity of the spinon-mediated tunnel spin current observed in the previous experiment [62], which is $10^{-4}$$10^{-3}$ times smaller than that of the 3D FM state, the signal intensity in the KSLs is expected to be sufficiently large to be observed experimentally. The signal in a typical 3D FM state is of the order of $1$ to $10~{}\mu\mathrm{V}$ in a typical setup [61]. Therefore, form these comparisons, the signal intensity in the KSLs is roughly estimated to be in the range of $10~{}\mathrm{nV}$$1~{}\mu\mathrm{V}$.
Table 1: Comparison of the tunnel spin current $\tilde{J}_{S}$ between the KSL and typical 3D FM states. The data for KSL are of $(X,L)=(6,10)$ and ${\bf B}\parallel c$ extracted from Fig. 4(c), while those for the FM states are computed using the $S=1/2$ Heisenberg model on a simple cubic lattice with $B/J=0.1$ based on the spin-wave theory (see Appendix A).
System State $T$ $\tilde{J}_{S}$
3D Heisenberg FM $0.1J$ 0.178
3D Heisenberg FM $0.5J$ 4.73
2D FM Kitaev KSL $0.05|K|$ 0.111
2D AFM Kitaev KSL $0.05|K|$0.500
@@ V Summary
In summary, we have studied the spin Seebeck effect in both the ferromagnetic and antiferromagnetic Kitaev spin liquids by using the real-time dynamics simulation, the perturbation analysis, and the linear spin-wave theory, based on the tunnel spin-current theory. There are two main discoveries. The first one is that the spin Seebeck effect is induced in the Kitaev spin liquid by the low-energy fractional Majorana quasiparticles that have no spin angular momentum. This is a qualitatively different phenomenon from the spin Seebeck effect in the one-dimensional quantum spin liquid induced by the fractional spinon excitations with nonzero spin angular momentum. The second discovery is that the spin current mediated by elementary excitations in quantum spin liquids can be positive or negative; namely, we found the sign change of the spin current in the Kitaev spin liquid between the ferromagnetic and antiferromagnetic Kitaev interactions. Our finding suggests the Majorana fermions contribute to the spin current with up- and down-spin-like nature in the ferromagnetic and antiferromagnetic Kitaev spin liquids, respectively. This finding refutes a possibility that the spin current carried by the fractional excitations is always negative in the quantum spin liquids.
Fractional excitations in quantum spin liquids are hard to identify by conventional experimental probes, since they behave very differently from the original spin degree of freedom and there is no conjugate field to directly excite them. Our finding suggests that the spin Seebeck effect would provide a strong complementary method and would be a powerful tool for investigating the nature of fractional excitations, in addition to the existing methods, such as the thermal Hall effect [33, 35, 34, 36, 40], the inelastic neutron scattering [98, 99, 100, 101, 95, 102, 103, 104], the Raman scattering [108, 109, 110], the electron spin resonance [105], and the nuclear magnetic resonance [106, 107].
Our results show that the Majorana quasiparticles in the Kitaev spin liquid contribute to the spin current, suggesting the possibility of generating and controlling them by a spin injection. It is highly desired to confirm this phenomenon experimentally by using the setup in Fig. 1(a) for both ferromagnetic and antiferromagetic Kitaev magnets; the former includes several iridium oxides and $\alpha$-RuCl3 [26, 27, 28], and the latter was theoretically predicted, e.g., for polar spin-orbit Mott insulators [113] and $f$-electron compounds [96, 97].
@@ Acknowledgements.
The authors would like to thank K. Fukui, H.-C. Jiang, K. Kobayashi, E. Saitoh, S. Trebst, and A. Tsukazaki for fruitful discussions. This work was supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP19H05825, No. JP20H00122, No. JP20H01830, No. JP20H01849, No. JP22H01179, No. JP22K03509, No. JP23H03818, No. JP23K22450, and No. JP24K00563, JSPS Grant-in-Aid for Scientific Research on Innovative Areas Grants No. JP19H05825, No. JP22H04480, No. JP22H05131, and No. JP23H04576, JST COI-NEXT Program Grant Number JPMJPF2221, and JST, CREST Grant No. JP-MJCR18T2 and No. JPMJCR24R5, Japan. This work was also supported by the National Natural Science Foundation of China (Grant No. 12150610462). Numerical calculations were performed using the facilities of the Supercomputer Center, The Institute for Solid State Physics, The University of Tokyo.
@@ Appendix A Spin current in the ferromagnetic Heisenberg model
In this Appendix, we show how to compute the tunnel spin current $\tilde{J}_{S}$ in Eq. (6) for the $S=1/2$ FM Heisenberg model [72, 62, 76] shown in Fig. 4(c). The Hamiltonian reads
@@
where the sum of $\langle j,j^{\prime}\rangle$ is taken for all the nearest-neighbor spins on the honeycomb lattice; we take $J=1$. Without loss of generality, we consider ${\bf B}\parallel z$, for which all the spins are fully polarized to the $z$ direction in the ground state. The following calculations are described using the model on a honeycomb lattice, but extensions to models on other lattices are straightforward. We show the results for a simple cubic lattice in Sec. IV.4.
We calculate $\tilde{J}_{S}$ for the fully-polarized (FP) state by using the linear spin-wave theory. By the Holstein-Primakoff transformation, $S^{+}_{j}=b^{\;}_{j}$, $S^{-}_{j}=b^{\dagger}_{j}$, and $S^{z}_{j}=\frac{1}{2}-n_{j}$ with $n_{j}=b^{\dagger}_{j}b^{\;}_{j}$, the Hamiltonian is written by the boson operators $b^{\;}_{j}$ and $b^{\dagger}_{j}$ as
@@
$\mathcal{H}^{\rm LSW}_{{\rm H}}=$ $-\sum_{\bf q}\left(g_{\bf q}b^{\dagger}_{{\bf q}\rm B}b_{{\bf q}\rm A}+{\rm h.c.}\right)$
@@
where $g_{\bf q}=(J/2)(e^{i{\bf q}\cdot{\bf a}_{1}}+e^{i{\bf q}\cdot{\bf a}_{2}}+1)$ with two translation vectors ${\bf a}_{1}$ and ${\bf a}_{2}$, and the index $\alpha$ denotes the sublattice $\mathrm{A}$ and $\mathrm{B}$. This Hamiltonian is easily diagonalized as
$\mathcal{H}^{\rm LSW}_{{\rm H}}=$ $\sum_{{\bf q}}\sum_{\sigma=\pm}\varepsilon_{{\bf q}\sigma}\eta^{\dagger}_{{\bf q}\sigma}\eta^{\;}_{{\bf q}\sigma},$ (20)
@@
$X_{+}(\omega)$ $=\frac{\delta}{2}\sum_{\ell}\frac{|\langle\ell|b^{\dagger}_{j_{0}}|{\rm FP}\rangle|^{2}}{(E_{\ell}-E_{\rm FP}-\omega)^{2}+\delta^{2}}$
$=\frac{\delta}{2N_{\rm u}}\sum_{{\bf q},\sigma}\frac{1}{(\varepsilon_{{\bf q}\sigma}-\omega)^{2}+\delta^{2}}$
@@ Appendix B Matrix elements of $s^{\pm(1)}$ and $s^{\pm(2)}_{1}$
where the sum runs over half of the first Brillouin zone with $q_{x}>0$ and $\alpha=\mathrm{A}$ or $\mathrm{B}$, Eq. (8) is rewritten as
$\mathcal{H}_{0}=$ ${\sum_{\bf q}}^{\prime}{\boldsymbol{\gamma}}^{\dagger}_{\bf q}\mathsf{H}_{\bf q}{\boldsymbol{\gamma}}^{\;}_{\bf q}$ (26)
@@
$\mathsf{U}_{\bf q}=\frac{1}{\sqrt{2}|f_{\bf q}|}\left(\begin{matrix}|f_{\bf q}|&if_{\bf q}\\ -if^{*}_{\bf q}&-|f_{\bf q}|\end{matrix}\right)$ (28)
@@
$\mathcal{H}_{0}=$ ${\sum_{\bf q}}^{\prime}\sum_{\sigma=\pm}\sigma|f_{\bf q}|\alpha^{\dagger}_{{\bf q}\sigma}\alpha^{\;}_{{\bf q}\sigma},$ (29)
where ${\boldsymbol{\alpha}}^{\dagger}_{\bf q}=(\alpha^{\dagger}_{{\bf q}+}~{}\alpha^{\dagger}_{{\bf q}-})$ and ${\boldsymbol{\alpha}}^{\dagger}_{\bf q}={\boldsymbol{\gamma}}^{\dagger}_{\bf q}\mathsf{U}_{\bf q}$.
@@
Note that the energy of $|\ell\rangle=|{\bf q}-,{\bf q}^{\prime}-\rangle$ and $|\ell^{\prime}\rangle=|{\bf q}+,{\bf q}^{\prime}+\rangle$ is the same, since $E_{\ell}=E_{\ell^{\prime}}=|f_{\bf q}|+|f_{{\bf q}^{\prime}}|$. For the particle-hole states, the matrix element is given by
@@
The matrix elements for $s^{-(2)}_{1}$ are obtained by replacing $A^{\gamma}$ by $(A^{\gamma})^{*}$. If the sign of $K$ is reversed, $f_{\bf q}$ becomes $-f_{\bf q}$, and then, $\langle{\bf q}\pm,{\bf q}^{\prime}\pm|s^{\pm(1)}|\emptyset\rangle$ and $\langle{\bf q}+,{\bf q}^{\prime}-|s_{1}^{\pm(2)}|\emptyset\rangle$ change their signs, while $\langle{\bf q}+,{\bf q}^{\prime}-|s^{\pm(1)}|\emptyset\rangle$ and $\langle{\bf q}\pm,{\bf q}^{\prime}\pm|s_{1}^{\pm(2)}|\emptyset\rangle$ are intact. Therefore, the sign reversal of $K$ leads to $\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle\to-\langle\emptyset|s^{\mp(2)}_{1}|\ell\rangle\langle\ell|s^{\pm(1)}|\emptyset\rangle$.